# 3D-Mem: 3D Scene Memory for Embodied Exploration and Reasoning

### v4

@@ Abstract
Constructing compact and informative 3D scene representations is essential for effective embodied exploration and reasoning, especially in complex environments over extended periods. Existing representations, such as object-centric 3D scene graphs, oversimplify spatial relationships by modeling scenes as isolated objects with restrictive textual relationships, making it difficult to address queries requiring nuanced spatial understanding. Moreover, these representations lack natural mechanisms for active exploration and memory management, hindering their application to lifelong autonomy. In this work, we propose 3D-Mem, a novel 3D scene memory framework for embodied agents. 3D-Mem employs informative multi-view images, termed Memory Snapshots, to represent the scene and capture rich visual information of explored regions. It further integrates frontier-based exploration by introducing Frontier Snapshots—glimpses of unexplored areas—enabling agents to make informed decisions by considering both known and potential new information. To support lifelong memory in active exploration settings, we present an incremental construction pipeline for 3D-Mem, as well as a memory retrieval technique for memory management. Experimental results on three benchmarks demonstrate that 3D-Mem significantly enhances agents’ exploration and reasoning capabilities in 3D environments, highlighting its potential for advancing applications in embodied AI.
@@ 1 Introduction
However, these scene representations exhibit significant limitations when employed as scene memory for embodied agents. Object-centric representations like 3D scene graphs tend to oversimplify 3D scenes by modeling them as individual objects and quantizing inter-object relationships into restrictive textual descriptions. This approach results in the loss of critical spatial information, making it difficult to answer questions that require an understanding of spatial relationships. As illustrated in the bottom-left example from Figure 1, when attempting to answer the question “Is there enough room to place a coffee table in front of the armchair?” using a 3D scene graph, the only spatial-related information we can access are the numerical 3D bounding boxes of the objects and the textual descriptions of the spatial relationships among them. It is challenging to measure unoccupied space or determine oriented spatial relationships such as “in front of” based solely on this information. On the other hand, dense 3D representations like point clouds or neural fields [41, 39, 10, 13] are computationally expensive and lack scalability as the scene grows during the agent’s exploration. Meanwhile, unlike vision-language models (VLMs) and large language models (LLMs) that effectively reason over images and text, current foundation models lack sufficient reasoning capabilities for dense 3D modalities due to limited training data. Additionally, neither type of existing scene representation can model unexplored regions, thus failing to support agents in active exploration, which restrains agents from leveraging their scene memory to expand their knowledge and achieve their goals for embodied tasks.
@@
To address these challenges, we introduce 3D-Mem, a more capable 3D scene memory for embodied agents. 3D-Mem employs a set of informative multi-view images, termed Memory Snapshots, to encompass the explored regions of a 3D scene. Each memory snapshot captures all objects visible in that snapshot along with their surroundings. Our intuition is that a snapshot image alone is sufficient to capture rich visual information of a region. For example, revisiting the challenge illustrated in Figure 1, a memory snapshot clearly shows that there is sufficient space “in front of” the armchair to place a coffee table. With the recent advancements in VLMs, these models can extract such spatial information from images [2, 23], much like humans do through intuitive observations.
@@
Building upon the concept of employing multi-view images as 3D scene memory, 3D-Mem integrates with frontier-based exploration frameworks and extends the concept of “frontiers” to Frontier Snapshots, which represent glimpses of unexplored regions in a 3D scene. In the scenario depicted in the top-left example from Figure 1, we can observe that the frontier snapshot corresponds to a corridor, which is the most probable path leading to the front door. By maintaining these frontier snapshots, the agent can choose either to complete tasks based on its accumulated knowledge or navigate to unexplored regions for new information, as illustrated in the embodied question-answering tasks in Figure 1. Meanwhile, representing both explored and unexplored regions through multi-view images allows us to better leverage the decision-making capability of VLMs, enhancing the agent’s ability to reason and plan effectively.
@@
Additionally, as agents in lifelong settings continuously explore and expand their knowledge, the 3D scene memory system must operate effectively when the memory grows large. To achieve this, 3D-Mem supports real-time incremental memory aggregation. To efficiently manage the ever-expanding scene memory, we introduce Prefiltering as an effective memory retrieval mechanism to select relevant memory during decision-making. This framework enables the agent to perform continuous exploration and navigation over extended periods without excessive computational burdens.
@@ 2 Related Work
3D Scene Representations. Recent works [25, 43, 33, 40] have focused on establishing dense 3D representations by grounding 2D representations captured by Vision-Language Foundation Models (e.g. CLIP [27], BLIP [20], SEEM [45]) into 3D scenes, which showcase impressive results on tasks such as open-vocabulary object segmentation and language-guided object grounding [9]. However, such representations are limited due to high resource consumption and the inability to support dynamic updates. Moreover, their scene features are limited to the semantic level and do not include spatial and multi-scale information. 3D scene graphs address these limitations by formulating the scene as a compact graph, where nodes represent objects, and edges encode inter-object relationships as textual descriptions [7, 1, 19, 37, 8], enabling real-time establishment and dynamic update for hierarchical scene representations [30, 38, 14]. While such object-centric representations have demonstrated effectiveness in various tasks, they remain constrained for oversimplifying inter-object relationships with rigid descriptions and missing mechanisms for active exploration and memory management. To tackle this challenge, our work leverages a set of informative memory snapshots to visually capture spatial and semantic relationships among objects, offering a more sophisticated understanding of the scene.
VLM for Exploration and Reasoning. Vision-Language Models (VLMs) have shown promising results in solving embodied exploration and reasoning tasks by leveraging commonsense reasoning and internet-scale knowledge. Existing exploration approaches can be divided into two categories. The former directly employs consecutive observations together with instructions as input, requiring the VLM to predict the next-step action [44] while the latter grounds the exploration target in the 3D scene through visual prompting, establishing a semantic map to guide the exploration process [22, 34, 29, 42]. However, both approaches are constrained by their memory representations. For the former, vanilla past observations can only serve as short-term memory. For the latter, their semantic maps are target-specific and cannot be generalized to future tasks. On the other hand, current reasoning approaches generally assume a fully observable scene as the input of the VLM, either represented with image observations [2] or 3D scene representations like point clouds [11], which makes them inapplicable in partially mapped environments. To address these limitations, our work introduces the first lifelong and target-agnostic scene memory that can be seamlessly integrated with VLM for further reasoning, stepping closer to the ultimate goal of lifelong autonomy.
@@ 3.1.2 Co-Visibility Clustering
3.2 3D-Mem with Frontier-based Dynamic Exploration
@@ 3.2.2 Incremental Construction of 3D-Mem
Object Update. As illustrated in Figure 2, at each time step $t$, the agent first captures $N$ egocentric views $\mathcal{I}^{obs}=\{I_{1}^{obs},I_{2}^{obs},...,I_{N}^{obs}\}$, which are used to extract the object set $\mathcal{O}$ and the frame candidate set $\mathcal{I}$ similar to Section 3.1. Specifically, a threshold “$max\_dist$” is set to ensure only objects within a certain distance from the agent are added to the scene graph, as the memory snapshot should only represent objects from a local area. It is important to note that the object set $\mathcal{O}$ detected in these egocentric views may contain both newly identified objects and those already present in the previous set $\mathcal{O}_{t-1}$. Subsequently, the full object set and frame candidate set are updated as $\mathcal{O}_{t}=\mathcal{O}_{t-1}\cup\mathcal{O}$ and $\mathcal{I}_{t}=\mathcal{I}_{t-1}\cup\mathcal{I}$, respectively.
@@ 4 Experiments
As a form of 3D scene memory, 3D-Mem stores rich and compact visual information, serving as an effective framework for a lifelong agent to explore and reason about a 3D scene. To comprehensively evaluate 3D-Mem, we begin with Active Embodied Question Answering (Section 4.1), where the scene is initially unknown. This assessment tests 3D-Mem’s overall performance in scenarios that require both embodied exploration and reasoning. Next, we examine 3D-Mem’s efficiency in representing 3D scene information through Episodic Memory Embodied Question Answering (Section 4.2). In this evaluation, the scene scan of the ground truth region is provided and no exploration is needed for question answering. Following this, we evaluate 3D-Mem on GOAT-Bench (Section 4.3), a multi-modal lifelong navigation benchmark, to demonstrate 3D-Mem’s effectiveness as a lifelong memory system. Finally, we conduct a series of ablation studies to determine key hyperparameter choices in Appendix 12.
@@ 4.1 Active Embodied Question Answering
Table 1: Experiments on A-EQA. “CG” denotes ConceptGraphs. Methods with * are reported from OpenEQA [23].
Benchmark. A-EQA consists of 557 questions drawn from 63 scenes in HM3D [28]. Due to resource limitations, our evaluation focuses on a subset of 184 questions, as mentioned in the OpenEQA benchmark [23]. The open-vocabulary and open-ended questions in A-EQA encompass diverse daily tasks such as object recognition, functional reasoning, and spatial understanding. For each question, an agent is initialized at a specific location and is required to explore the scene to gather the necessary information for answering the question.
Metrics. Following OpenEQA, we employ LLM-Match and LLM-Match SPL for quantitative evaluation. We first rate each predicted answer from 1 to 5 using GPT-4 to compare ground-truth and predicted answers. Given the predicted answers, LLM-Match, which measures the answer accuracy, is calculated as the average score for each question, mapped to a 20-100 scale. LLM-Match SPL, which measures the exploration efficiency, is then calculated by weighting the LLM-Match score by exploration path length. For the questions where the VLM Exploration methods failed to provide an answer, we ask GPT-4o to directly guess an answer without visual inputs, setting the SPL to 0.0.
@@
4.2 Episodic-Memory Embodied QA)
@@ 4.3 GOAT-Bench
Benchmark. GOAT-Bench is a multimodal lifelong navigation benchmark, where an agent is tasked with sequentially navigating to several objects in an unknown scene, with each target described by either a category name (e.g., microwave), a language description (e.g., the microwave on the kitchen cabinet near the fridge), or an image of the target object. Due to the large size of GOAT-Bench and the resource limitations, we assess a subset of the “Val Unseen” split, consisting of one exploration episode for each of the 36 scenes, totaling 278 navigation subtasks.
@@
Table 3: Experiments on GOAT-Bench. Evaluated on the “Val Unseen” split. “CG” denotes ConceptGraphs. Methods denoted by * are from GOAT-Bench.
@@ 5 Conclusion
We present 3D-Mem, a 3D scene memory framework that uses a set of informative multi-view snapshot images to cover the scene and store robust visual information. With the integration of the frontier-based exploration framework, 3D-Mem allows the agent to either leverage the memory of explored regions to solve tasks or explore the scene to expand its knowledge. With its incremental construction and efficient memory retrieval mechanism, 3D-Mem serves as an effective memory system for lifelong agents. Extensive experiments demonstrate the significant advantages of 3D-Mem over traditional scene representations.
@@ 6.1 Detailed Experiment Results
Table 4: Performance on A-EQA by Question Categories. For each question categories, there are two columns. The first column stands for the LLM-Match Score, while the second column represents the SPL score. “CG” denotes ConceptGraphs. Methods with * are reported from OpenEQA [23]. Columns represent different category of questions in the dataset.
@@
Table 5: Performance on GOAT-Bench by Question Modalities. Evaluated on the “Val Unseen” split. “CG” denotes ConceptGraphs. Methods denoted by * are from GOAT-Bench.

### v5

@@ Abstract
Constructing compact and informative 3D scene representations is essential for effective embodied exploration and reasoning, especially in complex environments over extended periods. Existing representations, such as object-centric 3D scene graphs, oversimplify spatial relationships by modeling scenes as isolated objects with restrictive textual relationships, making it difficult to address queries requiring nuanced spatial understanding. Moreover, these representations lack natural mechanisms for active exploration and memory management, hindering their application to lifelong autonomy. In this work, we propose 3D-Mem, a novel 3D scene memory framework for embodied agents. 3D-Mem employs informative multi-view images, termed Memory Snapshots, to capture rich visual information of explored regions. It further integrates frontier-based exploration by introducing Frontier Snapshots—glimpses of unexplored areas—enabling agents to make decisions by considering both known and potential new information. To support lifelong memory in active exploration settings, we present an incremental construction pipeline for 3D-Mem, as well as a memory retrieval technique for memory management. Experimental results on three benchmarks demonstrate that 3D-Mem significantly enhances agents’ exploration and reasoning capabilities in 3D environments, highlighting its potential for advancing applications in embodied AI.
@@ 1 Introduction
However, these scene representations exhibit significant limitations when employed as scene memory for embodied agents. Object-centric representations like 3D scene graphs tend to oversimplify 3D scenes by modeling them as individual objects and quantizing inter-object relationships into restrictive textual descriptions. This approach results in the loss of critical spatial information, making it difficult to answer questions that require an understanding of spatial relationships. As illustrated in the bottom-left example from Figure 1, when attempting to answer the question “Is there enough room to place a coffee table in front of the armchair?" using a 3D scene graph, the only spatial-related information we can access are the numerical 3D bounding boxes of the objects and the textual descriptions of the spatial relationships among them. It is challenging to measure unoccupied space or determine oriented spatial relationships such as “in front of" based solely on this information. On the other hand, dense 3D representations like point clouds or neural fields [40, 38, 10, 12] are computationally expensive and lack scalability as the scene grows during the agent’s exploration. Meanwhile, unlike vision-language models (VLMs) and large language models (LLMs) that effectively reason over images and text, current foundation models lack sufficient reasoning capabilities for dense 3D modalities due to limited training data. Additionally, neither type of existing scene representation can model unexplored regions, thus failing to support agents in active exploration, which restrains agents from leveraging their scene memory to expand their knowledge and achieve their goals for embodied tasks.
@@
To address these challenges, we introduce 3D-Mem, a more capable 3D scene memory for embodied agents. 3D-Mem employs a set of informative multi-view images, termed Memory Snapshots, to encompass the explored regions of a 3D scene. Each memory snapshot captures all objects visible in that snapshot along with their surroundings. Our intuition is that a snapshot image alone is sufficient to capture rich visual information of a region. For example, revisiting the challenge illustrated in Figure 1, a memory snapshot clearly shows that there is sufficient space “in front of" the armchair to place a coffee table. With the recent advancements in VLMs, these models can extract such spatial information from images [2, 22], much like humans do through intuitive observations.
@@
Building upon the concept of employing multi-view images as 3D scene memory, 3D-Mem integrates with frontier-based exploration frameworks and extends the concept of “frontiers" to Frontier Snapshots, which represent glimpses of unexplored regions in a 3D scene. In the scenario depicted in the top-left example from Figure 1, we can observe that the frontier snapshot corresponds to a corridor, which is the most probable path leading to the front door. By maintaining these frontier snapshots, the agent can choose either to complete tasks based on its accumulated knowledge or navigate to unexplored regions for new information, as illustrated in the embodied question-answering tasks in Figure 1. Meanwhile, representing both explored and unexplored regions through multi-view images allows us to better leverage the decision-making capability of VLMs, enhancing the agent’s ability to reason and plan effectively.
@@
Additionally, as agents in lifelong settings continuously explore and expand their knowledge, the 3D scene memory system must operate effectively as the memory grows. To achieve this, 3D-Mem supports real-time incremental memory aggregation. To efficiently manage the ever-expanding scene memory, we introduce Prefiltering as an effective memory retrieval mechanism to select relevant memory during decision-making. This framework enables the agent to perform continuous exploration and navigation over extended periods without excessive computational burdens.
@@ 2 Related Work
3D Scene Representations. Recent works [24, 42, 32, 39] have focused on establishing dense 3D representations by grounding 2D representations captured by Vision-Language Foundation Models (e.g. CLIP [26], BLIP [19], SEEM [44]) into 3D scenes, which showcase impressive results on tasks such as open-vocabulary object segmentation and language-guided object grounding [9]. However, such representations are limited due to high resource consumption and the inability to support dynamic updates. 3D scene graphs address these limitations by formulating the scene as a compact graph, where nodes represent objects, and edges encode inter-object relationships as textual descriptions [7, 1, 18, 36, 8], enabling real-time establishment and dynamic update for hierarchical scene representations [29, 37, 13]. While such object-centric representations have demonstrated effectiveness in various tasks, they remain constrained for oversimplifying inter-object relationships with restrictive text descriptions. To tackle this challenge, our work leverages a set of informative memory snapshots to visually capture spatial and semantic relationships among objects, offering a more sophisticated understanding of the scene.
VLM for Exploration and Reasoning. Vision-Language Models (VLMs) have shown promising results in solving embodied exploration and reasoning tasks by leveraging commonsense reasoning and internet-scale knowledge. Existing exploration approaches can be divided into two categories. The former directly employs consecutive observations together with instructions as input, requiring the VLM to predict the next-step action [43] while the latter grounds the exploration target in the 3D scene through visual prompting, establishing a semantic map to guide the exploration process [21, 33, 28, 41]. However, both approaches are constrained by their memory representations. For the former, vanilla past observations can only serve as short-term memory. For the latter, their semantic maps are target-specific and cannot be generalized to future tasks. To address these limitations, our work introduces the first lifelong and target-agnostic scene memory that can be seamlessly integrated with VLM for further reasoning, stepping closer to the ultimate goal of lifelong autonomy.
Topological Mapping. Besides the 3D scene representations mentioned above, prior 2D topological mapping methods like TSGM [17] and RoboHop  [6] provide valuable context. They construct navigation graphs from images and objects but do not fully capture all objects or their interrelationships. In contrast, 3D-Mem clusters co-visible objects to represent inter-object spatial relationships, enabling tasks such as embodied Q&A that extend beyond route planning. Moreover, its efficient memory retrieval mechanism scales to lifelong exploration, leveraging VLMs for advanced reasoning and decision-making.
@@ 3.1.2 Co-Visibility Clustering
3.2 3D-Mem with Frontier-based Exploration
@@ 3.2.2 Incremental Construction of 3D-Mem
Object Update. As illustrated in Figure 2, at each time step $t$, the agent first captures $N$ egocentric views $\mathcal{I}^{obs}=\{I_{1}^{obs},I_{2}^{obs},...,I_{N}^{obs}\}$, which are used to extract the object set $\mathcal{O}$ and the frame candidate set $\mathcal{I}$ similar to Section 3.1. Specifically, a threshold “$max\_dist$" is set to ensure only objects within a certain distance from the agent are added to the scene graph, as the memory snapshot should only represent objects from a local area. It is important to note that the object set $\mathcal{O}$ detected in these egocentric views may contain both newly identified objects and those already present in the previous set $\mathcal{O}_{t-1}$. Subsequently, the full object set and frame candidate set are updated as $\mathcal{O}_{t}=\mathcal{O}_{t-1}\cup\mathcal{O}$ and $\mathcal{I}_{t}=\mathcal{I}_{t-1}\cup\mathcal{I}$, respectively.
@@ 4 Experiments
As a form of 3D scene memory, 3D-Mem stores rich and compact visual information, serving as an effective framework for a lifelong agent to explore and reason about a 3D scene. To comprehensively evaluate 3D-Mem, we begin with Active Embodied Question Answering (Section 4.1), where the scene is initially unknown. This assessment tests 3D-Mem’s overall performance in scenarios that require both embodied exploration and reasoning. Next, we examine 3D-Mem’s efficiency in representing 3D scene information through Episodic Memory Embodied Question Answering (Section 4.2). In this evaluation, the scene scan of the ground truth region is provided and no exploration is needed for question answering. Following this, we evaluate 3D-Mem on GOAT-Bench (Section 4.3), a multi-modal lifelong navigation benchmark, to demonstrate 3D-Mem’s effectiveness as a lifelong memory system. Finally, we conduct a series of ablation studies to determine key components and hyperparameter choices in Appendix 13.
@@ 4.1 Active Embodied Question Answering
[…]
Benchmark. A-EQA consists of 557 questions drawn from 63 scenes in HM3D [27]. Due to resource limitations, our evaluation focuses on a subset of 184 questions, as mentioned in the OpenEQA benchmark [22]. We also include the evaluation of 3D-Mem on the full set in Appendix 6. The open-vocabulary and open-ended questions in A-EQA encompass diverse daily tasks such as object recognition, functional reasoning, and spatial understanding. For each question, an agent is initialized at a specific location and is required to explore the scene to gather the necessary information for answering the question.
Metrics. Following OpenEQA, we employ LLM-Match and LLM-Match SPL for quantitative evaluation. We first rate each predicted answer from 1 to 5 using GPT-4 to compare ground-truth and predicted answers. Given the predicted answers, LLM-Match, which measures the answer accuracy, is calculated as the average score for each question, mapped to a 0-100 scale. LLM-Match SPL, which measures the exploration efficiency, is then calculated by weighting the LLM-Match score by exploration path length. For the questions where the VLM Exploration methods failed to provide an answer, we ask GPT-4o to directly guess an answer without visual inputs, setting the SPL to 0.0.
@@
4.2 Episodic-Memory Embodied Q&A
@@ 4.3 GOAT-Bench
Benchmark. GOAT-Bench is a multimodal lifelong navigation benchmark, where an agent is tasked with sequentially navigating to several objects in an unknown scene, with each target described by either a category name (e.g., microwave), a language description (e.g., the microwave on the kitchen cabinet near the fridge), or an image of the target object. Due to the large size of GOAT-Bench and the resource limitations, we assess a 1/10-size subset of the “Val Unseen” split, consisting of one exploration episode for each of the 36 scenes, totaling 278 navigation subtasks. For reference, we also include the evaluation of 3D-Mem on the full test set in Appendix 6.
@@
Table 3: Experiments on GOAT-Bench. Evaluated on the “Val Unseen" split. “CG” denotes ConceptGraphs. Methods denoted by * are from GOAT-Bench, and with $\dagger$ are evaluated on the subset.
@@ 5 Conclusion
We present 3D-Mem, a 3D scene memory framework that uses a set of informative multi-view snapshot images to store robust visual information of a 3D scene. With the integration of the frontier-based exploration framework, 3D-Mem allows the agent to either leverage the memory of explored regions to solve tasks or further explore the scene to expand its knowledge. With its incremental construction and efficient memory retrieval mechanism, 3D-Mem serves as an effective memory system for lifelong agents. Extensive experiments demonstrate the significant advantages of 3D-Mem over traditional scene representations.
@@
6 Full-Set Evaluation
Following the common practice and due to resource limitations, we only evaluate baselines and our method on a subset of A-EQA and GOAT-Bench in our main paper. For reference, we also evaluate 3D-Mem on the complete benchmarks, as shown in the following table.
Whole Set Subset
GOAT-Bench Success Rate $\uparrow$ SPL $\uparrow$ Success Rate $\uparrow$ SPL $\uparrow$
3D-Mem (Ours) 62.9 44.7 69.1 48.9
Whole Set Subset
A-EQA LLM-Match $\uparrow$ LLM-Match SPL $\uparrow$ LLM-Match $\uparrow$ LLM-Match SPL $\uparrow$
3D-Mem (ours) 53.3 38.0 52.6 42.0
@@ 7.1 Detailed Experiment Results
Table 4: Performance on A-EQA by Question Categories. For each question categories, there are two columns. The first column stands for the LLM-Match Score, while the second column represents the SPL score. “CG" denotes ConceptGraphs. Methods with * are reported from OpenEQA [22]. Columns represent different category of questions in the dataset.
@@
Table 5: Performance on GOAT-Bench by Question Modalities. Evaluated on the “Val Unseen" split. “CG” denotes ConceptGraphs. Methods denoted by * are from GOAT-Bench.
@@ 7.2 Decision Frequency
7.3 Limitations
We acknowledge several key constraints of 3D-Mem: (1) Similar to most 3D scene representations, 3D-Mem is designed for static environments and is not robust to moving objects. (2) The performance of 3D-Mem depends on object detection results and VLM reasoning capabilities. (3) The precise location of the agent is required to accurately locate the objects in the scene and construct scene memory, which could be challenging to acquire after long-time exploration. (4) The latency of the whole pipeline during embodied QA is still noticeable. We time the detailed latency of each component of our pipeline in Table 6. We also argue that, although 2D-3D lifting (including SLAM and object detection) and Prefiltering consume considerable time, in real-world scenarios, they run concurrently during navigation, and caching Prefiltering results minimizes their impact on throughput. The primary bottleneck is VLM inference, which is caused by both heavy model computation and network latency. This can be mitigated by running the model locally or optimizing VLMs with techniques such as model quantization. Additionally, 3D-Mem requires VLM inference only after each high-level step, thereby reducing the VLM query frequency.
Component 2D-3D Lift Clustering Prefiltering VLM Inference
A-EQA 2.43 0.04 1.12 3.34
GOAT-Bench 2.79 0.09 1.35 3.58
Table 6: Time cost of each component on A-EQA, evaluated in seconds.
@@ 13.1 Ablation on Hyperparameter Choices
13.2 Ablation on Pipeline Components
Ablation study on Prefiltering is infeasible because directly querying the VLM would exceed its context length. However, we conduct an ablation on Frontier Snapshots by always navigating to the nearest frontier when Memory Snapshots cannot provide the answer, rather than choosing a frontier via VLM. As shown in Table 7 (SnapMem w/o FS), performance declines on both A-EQA and GOAT-Bench, though the drop is smaller on GOAT-Bench. This is likely due to the lifelong setting of GOAT-Bench, where the agent tends to rely on its memory once the scene is mostly explored. Additional experiments removing both Frontier Snapshots and memory maintenance (SnapMem w/o FS & Mem) confirm this pattern.
A-EQA GOAT-Bench
Method LLM-Match $\uparrow$ LLM-Match SPL $\uparrow$ Success Rate $\uparrow$ SPL $\uparrow$
SnapMem w/o FS & Mem - - 57.2 33.2
SnapMem w/o FS 49.3 31.0 63.7 46.8
SnapMem 52.6 42.0 69.1 48.9
Table 7: Ablation study of Frontier Snapshot on A-EQA and GOAT-Bench. FS denotes "Frontier Snapshots".